Please someone help me! I need to rewrite this function in turning point form
y=x^+4x+3

Answer by stanbon(75887)

rewrite this function in turning point form
y=x^2+4x+3
The form is y-k=(x-h)^2 where (h,k) is the vertex or turning point.
Rewrite as:
y-3 = x^2+4x
Complete the square on the right side and keep the equation balanced:
y-3+? = x^2+4x+?
y--3+4 = x^2+4x+4
y+1 = (x+2)^2
Therefore:
The vertex or "turning point" is at x-2 , y=-1
Cheers,
Stan H.